Presentation is loading. Please wait.

Presentation is loading. Please wait.

Tinker AFB – F100-200, F14 jet engine test MTT, TBO 3/29/07 – 4/6/07.

Similar presentations

Presentation on theme: "Tinker AFB – F100-200, F14 jet engine test MTT, TBO 3/29/07 – 4/6/07."— Presentation transcript:

1 Tinker AFB – F100-200, F14 jet engine test MTT, TBO 3/29/07 – 4/6/07

2 Log of SI and Airbeam March 30, 2007 MCP = 2220 V; IE = 8.5e-6 March 31, 2007 MS AB = 9e5; PTOF AB = 1.2e6 ;SI = 18; MS AB/SI = 5e4 April 2, 2007 MORNING MS AB = 7e5; PTOF AB = 8e5; SI = 14; MS AB/SI = 5e4 April 2, 2007 LATE MORNINGS MS AB = 6.4e5; PTOF AB 7e5; SI = 13; MS AB/SI = 4.9e4 April 2, 2007 11:47 AM MCP = 2260 V from 2220 V April 2, 2007 11:47am, MS AB = 8e5; PTOF AB = 6e5; SI = 17; MS AB/SI = 4.7e4 WORRIED April 2, 2007 12:00pm, MS AB = 8.5e5; PTOF AB = 8.5e5; SI = 17; MS AB/SI = 5e4. THAT’S BETTER! April 3, 2007 8:45am, MS AB = 8e5; PTOF AB = 9e5; SI = 15; MS AB/SI = 5.3e4 April 3, 2007 1:00pm MS AB = 8.5e5; PTOF = 9.5e5; SI = 14; MS AB/SI = 6e4 (probably still seeing N2 enriched gas when this was measured) April 3, 2007 1:05PM MCP = 2280 V from 2260 V April 3, 2007 1:05pm SI = 14.5 MS AB = 7.5e5; PTOF AB = 8.5e5; SI = 14.5 MS AB/SI = 5.1e4 April 3, 2007 1:50pm SI = 14.5, MS AB = 7.9e5; PTOF AB = 9e5; MS AB/SI = 5.4e4 April 3, 2007 6:48pm SI = 13.5 MS AB = 8e5; PTOF AB = 9e5 Not too bad April 4, 2007 1:10pm SI = 14, MS AB = 8e5; PTOF AB = 8.5e5 MS AB/SI = 5.7e4

3 3/29/2007 Arrive at Tinker AFB Turn on pumps, oven, filament, MCP Performance is reasonable, though there is a substantial electric background The following slides document this noise issue

4 This is with the MCP fully connected

5 This is with the MCP connected but without the over-voltage protector

6 This is with a 50 ohm terminator

7 WTF! AMS meltdown Checked lots of things 1) external connections 2) chopper 3) Acq. Board connections 4) cold start…..everything checks out OK….error message “Is the chopper on?” Is it a timing issue with the board? After about 1 hr, bang – the signal came back. We have no satisfactory answers at this point…

8 PTOF looks OK, MS is a little strange…why does the signal go down?

9 Looks OK now

10 3/30/07 Calibrations Flow cal Size cal tuning

11 Looks OK this morning – noise may be a bit better? Zeros at a threshold of 5 instead of 6






17 070330_tinker_no_amp_noK, Heater bias = 37.4 V

18 And here’s what we get after about 1 hr of tuning, Heater bias = 37.4 V

19 Don’t pay attention to the SI for this screen, the noise though is very good!

20 This is the bitwise for IE_cal

21 This is the 10_sec GA bitwise So the SI is a bit different for the two menus

22 AMS run #’s 18375-18382 – with TPS settings shown above, looks very nice

23 3/31/07 1)Checking normal operation of 10-sec GA mode 2)Comparing different instruments







30 Found an error in the defaults morning Of 3/31/07

31 Re-checking the baseline in GA-10sec – looks very good

32 PTOF looks OK, what’s the deal with the MS AB? Is the servo moving faster again? Very strange…I will stick with the 300 msec b/c it’s working.

33 Screen shot during normal operation, Sat Morning at Tinker. So far so good. Good diff/closed air beams, reasonably low potassium. Good noise level. Saved a few runs 18388-18428 these are filter samples

34 Here’s the diagnostics from the filter samples run on the morning of 3/31/07


36 This may be an aggressive baseline setting – peaks at 280 are baseline away a bit. With these settings, I was trying to finesse the NH4 (15,16,17) peaks to see if I could keep them positive.

37 Mass spec looks OK Mass loadings look alright – NH4 is negative, org is noisy PTOF tof is strange – NH4 has a peak, org goes way neg PTOF size is strange – org looks like it will integrate to a negative # I guess the DC markers will need attention… St Dev x 3 => Org = 0.086 ug/m^3 (about 100 ng/m^3) SO4 = 0.015 ug/m^3 NO3 = 0.024 ug/m^3 Chl = 0.012 ug/m^3 (nominally!) NH4 = 0.057 ug/m^3

38 Org PTOF Why does the org ptof go negative? 17,18,41,44 all go negative at small tof 44 is CO2 – the truck had about 1000 ppm of CO2 during runs so that’s why it’s got problems 16,30,33 have both a negative and positive going peak at low tof 14,28,29,32,33,40 all have positive peaks 14, 28,32,and 40 are all primary air peaks 29 and 33 are isotopic peaks [O(16)-O(17) and N(15)-N(14)] So, we either have a baseline problem or we have a frag list problem I removed all of the offending masses from frag_organic but there was no difference with the ptof… Looking at the mass spec 15,16,20, 41, and 44 have organic masses Not worried about 44 for now b/c CO2 was so high. 15,16 are NH4/H2O problems. RH was different than frag list thinks it should be Tried using the air beam correction – no difference that I could see I tried to fine tune the squirrel baseline, the next slide shows what I got to

39 Here’s a fairly wide view, in general things look OK, but we need to take a closer look to know for sure. The next 3 slides look at 10-20, 20-40, and 90- 100 – the 3 interesting sections of this plot

40 OK, here’s the water/ammonium peaks – these look very nice

41 OK, here’s the air peaks – once again they look nice. It looks like our resolution is not sufficient to wipe out the ringing either in 28 or 32.

42 And here are the peaks in the 90’s – I’m not sure what these are…C7H7 I guess. At any rate, we wipe out a ringing peak in mass 91 and everything looks good

43 Filter Sample Redux IE and Airbeam look OK Mass loadings look reasonable To do list – 1) e-mail DS about dynamic CO2 correction to mass 44, 2) thermocouple calibrations for J and K type, 3) e-mail JK (cc JJ), tell him thanks looks good (4) email Donna that “Air” PTOF does not work in Squirrel (v1.36) 5) download Igor updater onto laptop, mem stick

44 April 1, 2007 Ran morning ambient samples and afternoon/morning filter samples In the afternoon runs, TBO had a great idea to change the data delay from 250 usec to 300 usec – nothing else was changed. This small modification got the organic PTOF to look just right. The next two slides are screen shots of the new timing menu for GA 10-sec and the signal vs TOF for the afternoon filter samples Also, noticed that the m/z calibration was not so good in Squirrel – take a look at that first thing Monday April 2

45 Compare this to slide #24 from March 30, 2007

46 Compare this to the previous two days – now that makes sense!

47 April 2, 2007 Tinker AFB Day 1 – Run #1 Filter samples A @ 7:51am, #18938; B 8:26am, #18997; C 8:45am, #19007; D 8:52am, #19018 As of 9:00am, the SI has decreased from about 18 on Sat to 14, AB beams had lost about 20% of there values as well PTOF, which had looked strange on Sat, but better after switching Data Delay from 250 usec to 300 usec, looks strange again. Increased chopper speed to 99.7 Hz (the set point is 100 Hz, and the speed had dropped to about 97 Hz) – PTOF may look a bit more normal. Increased data delay to 350 usec, but this had no effect. Try 320 usec.

48 Single ion decreased a bit…from 18 to 12? Try again after a few minutes of warm up time – it looks like the signal decreased in m/z cal as well; air beams are lower too

49 OK, this is a bit better – I forgot to set the pk discrimination to 4 in the previous slide, still we have gone from 18 to about 14.

50 400 microsec looks best – run with this setting 375 microsec

51 Just checking the servo – it looks good

52 Let’s try some different parameters What do you know, that seemed to work! The ptof is pretty noisy, but this is a small slice of data (9 runs) Let’s switch it back to 2000 nm by taking 350 spectra/chopper and see if this problem comes back

53 And there you go! The problem came back Let’s try to get rid of it again

54 Yikes! These last few slides document my efforts to “solve” the organic ptof problem… I was varying data delay and the # of extractions/chopper to find the sweet spot Suddenly, the PTOF AB got noisy. It would bounce from 8e5 to 1e5 and then back. And then, the PTOF AB started going negative, e.g. -8e5 and such. Called TBO…turned off computer, unplugged it (unplug fan first), pressed power switch, waited 1 minute or so, plugged computer back in, turned power back on and re-booted…had to toggle PTOF and re-baseline, but otherwise everything was cool Also, the PTOF AB, which had been 1e6 and then dropped to 8e5, was back to 1e6. Problem solved! But WTF is the root cause? It seems to be cross-talk between the DAQ board and software. What brings it on? This time, possibly switching the DAQ parameters in fairly rapid succession (or into a bad regime) may have led to the problems. One time, it looked like BP’s cell phone being turned off led to the same problem. This is definitely something to be aware of and to monitor.

55 Stay with what works for now! Why are the first few ORG points way negative? After about 4-5 runs it looks OK, but what is the beginning transient? It is NOT a problem with the MCP or with opening the valve. WTF is it?

56 DAY 1 Quick Notes MCP at 2260 V SI before and after about 15 – great stuff! Re-check tomorrow. Re-calibrate on Wednesday? Ran PTOF in sub-optimal mode Lots of organic at early time points – was the line dirty. Organic settled down with time, but there was still a measurable quantity at high powers. At low powers, there was much less AMS organic. CPC and SMPS looked good. MAAP looked very noisy! Can we replace the valve with a manifold of critical orifices? Send data to Tim. Closed valve, turned down MCP. Onto Squirrel!

57 Day 2 April 3, 2007 Plan – run dilution tests first with unheated (Run #2) and then heated lines (Run #3)

58 Check baseline and SI – no change, very good

59 PTOF At 8:35am, I tried to increase the # of spectra per chopper from 350 to 395. This increased the size range from 2000 nm to 3000 nm. The PTOF air beam immediately went negative! After one rotation, I quit and changed # of spectra back to 350. This seemed to correct the problem without requiring that I restart the computer/DAQ board. Saved data in Day2/prep folder – looked like normal in Squirrel

60 Runs 2 and 3 Checked SI = 15; MS AB = 8e5 @ 9:00am Run 2 begins 10:10am/#20461; ends 12:55pm/#21157 Checked SI = 14; MS AB = 8.4e5 Increased MCP to 2280 V Cleaned orifice @ 13:26 – checked flow Run #3 begins with filter samples at 13:50/#21160

61 Bitwise after Run#3 at 18:47 – we had been sitting at about 14.5 – not too bad Electronic noise is still reasonable

62 Donna/Tim Evaluation of Data Overall looks good Only problem is that 15-18 are lost probably due to high water background Ran filtered N2 through the AMS with the oven at 750 C overnight. DON’T FORGET TO TURN OVEN DOWN BEFORE RUNNING!

63 Day 3 – No Runs Preliminary look at 15-18 looked better Re-calibrated GA with BFSP gave a very small PTOF AB – ran BFSP alone and added GA MS/PTOF data later New IE = 8.19e-7, RIE = 2.02 (with MCP at 2280 V SI = 14)

64 MCP = 2280 V Noise looks awesome!

65 This is from the PTOF/MS file NO3/NH4 is about 7:1, it should be about 4:1




Download ppt "Tinker AFB – F100-200, F14 jet engine test MTT, TBO 3/29/07 – 4/6/07."

Similar presentations

Ads by Google